OpenStudy (anonymous):

log(x^9 + 8) to base e i need full steps

OpenStudy (anonymous):

the change of base formula when converting from base a to base b is\[\log_{a} x=\frac{\log_{b}x}{\log_{b}a} \]

OpenStudy (anonymous):

In your case you have a common log (base 10) that you want to change to base e (natural log) so relative to the formula a=10 and b=e:\[\log(x^9+8)=\frac{\ln (x^9+8)}{\ln 10}\]note that ln is a shorthand notation for "log to the base e"

OpenStudy (anonymous):

\[\log_e x=\ln x\]